OpenStudy (anonymous):

Factor the expression below. 9x2 - 64 Which of the following binomials is a factor of the expression?

OpenStudy (anonymous):

hint.. \[\LARGE 9=3^2\] and \[\LARGE 64=8^2\] now use this formula: \[\LARGE a^2-b^2=(a-b)(a+b)\] ...

OpenStudy (anonymous):

can you do it ?

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

What he's saying is you must use the formula so if you set a= 9 and b = 64 you need to square them right? \[3^{2}= 9\] and \[8^{2}= 64\] then simply plug into the formula to get \[3^{2}-8^{2}\] = \[(3-9)(3+9)\]

OpenStudy (anonymous):

A. 3x - 16 B. 3x - 8 C. 3x - 4 D. 3x - 32

OpenStudy (anonymous):

good example @Janex

OpenStudy (anonymous):

... unless there should be(3-8) ... anyway .. your answer is: \[\LARGE (3x)^2-8^2=(3x-8)(3x+8)\]

OpenStudy (anonymous):

oh that sould be 8 thx @kreshnik its = (3-8)(3+8)

OpenStudy (anonymous):

:) typing mistake. :P